# Balance Correlations, Agentic Zeros, and Networks: The Structure of 192 Years of War and Peace

### v1

@@ 4 Discussion and Extensions
This paper deals with signed graphs, where edges between nation states are undirected symmetric ties. One of Holland and Leinhardt’s major advances in this line of work was to recognize that networks are often, perhaps most often, composed of directed ties. The work presented here can be extended to directed ties, also, leading to new specific predictions and explorations of possible balance statements and commensurate balance correlations to assess them.
@@
However, it should be noted that this increases substantially the number of possible balance statements. Instead of one relation for each pair, there is both an in-tie and an out-tie that would have to be specified; thus, for each pair we must specify two relations, not just one. For example, we could describe a case where Ego sent a positive tie to Alter, but Alter does not respond with any sentiment (i.e., send a z-tie back). Perhaps Ego sends a negative tie to X, and X reciprocates with a non-tie. Further, suppose Alter sends a negative tie to X and X responds with a similar negative tie back to Alter. This configuration would be described as <pz . nz nn>. Since there are six directed ties in any labeled triple, then the number of possible directed triad signatures is $7293^{6}=$ 729, not simply $27$ as described in this paper. An interesting extension of our approach to digraphs is that of these $729$ signatures, $702$ have an indistinguishable twin and $27$ do not (see the proof of the theorem 2.1 in the appendix). We expound on these structures, their balance statements, and their empirical results in a subsequent paper on balance correlations in directed graphs.
5 Conclusion
@@
Our intent here is to propose a new approach to the theorizing and study of balance processes in network contexts. By formally incorporating neutral or zero ties into the analysis, we extend balance theory in a significant way that better matches the real world of actors in a systemic relational structure. And, in an important way, we return to Heider’s original thinking by focusing on Ego’s choice of how to relate to a specific Alter when confronted with the concommitant relations with one or many third parties. We present a novel nomenclature, the triad signature, that assists in identifying and theorizing about balanceimbalance conditions.
@@
Finally, we propose a simple, intuitive measure, the balance correlation, that directly reflects the extent to which Heider’s predictions (and its extentions in this paper) are supported by the evidence in an observed network. This measure captures the difference in probabilities of Ego’s choice of a tie with Alter, conditioned on whether the specific balance-theoretic 2-path configuration exists or does not exist. We also propose a reasonable null test of this statistical measure and apply it to a well-known data set employed by political scientists in international affairs, the Correlates of War data.
@@
Second, we find that the predictions around unbalanced states are not confirmed for the most part and occasionally even counter to Heider’s original contention. For example, the theoretically unbalanced <n.pp> correlation hovers around zero for most of the two centuries of data. And a clear example of a result that disconfirms an unbalanced prediction is given in Figure 6, where the theoretically unbalanced <p.pn> signature is consistently positive in the post WWII era.
Thirdly, in a related general finding, it appears that the z-ties (neutral ties) play a larger role in establishing balance, or at least avoiding unbalanced states, than negative hostile ties over substantial time periods. For example, in the case of where the Ego  Alter dyad is embedded in many <pp> 2-path relations (see Figure 5), there is do discernable evidence that Ego differentially avoids negative ties (<n.pp>), but there is substantial and significant suppression or avoidance of neutral ties (<z.pp>).
In sum, it is our contention that balance correlations provide a simple, intuitive way to assess the extent to which specific balance predictions might be true on the average across an entire network. Furthermore, by incorporating agentic zeroes into the exploration of balance conditions, we may uncover new properties of social order that have been prevalent, even dominant, but hidden from the researcher’s purvue or agenda. We hope this new approach opens up avenues of exploration that have been heretofore ignored.
Appendices
Appendix A Proof of theorem 2.1
The claim in Theorem 2.1 is based on the fact that if the indices of the two variables in the balance correlation are permuted in the same way, the correlation does not change due to associative and commutative properties of the addition of the cross-products in the correlation function. Transposing matrices is one unique permutation of the elements in a matrix.
Definition A.1.
@@
Proof of lemma A.1.
Note that $M_{q_{ji}}=M_{q_{ij}}^{{}^{\prime}}$. By contradiction, assume for any $ij$-th element in a nugget, the element-wise product of matrices, $M_{r_{ij}}\times M_{s_{ji}}=1$ and $M_{r_{ji}}\times M_{s_{ij}}=1$. Since, $M_{q_{ij}}\in\{0,1\}$, this implies $M_{r_{ij}}=M_{s_{ji}}=1$ and $M_{r_{ji}}=M_{s_{ij}}=1$, so $M_{r_{ij}}+M_{s_{ij}}=2$. However, if $r\neq s$ this contradicts that the sum over all matrices results in a matrix $\mathfrak{L}$ (see definition A.1). If and only if $r=s$ the assumption holds, as the summation visits each matrix $q$ in $\mathfrak{M}$ only once, which implies there are exactly $m$ symmetric nuggets; those of a matrix and it’s transpose. ∎
Definition A.3.
@@
$\rho(U_{k},U_{l}U_{m})=\frac{cov(U_{k},U_{l}U_{m})}{\sqrt{var(U_{k})var(U_{l}U_{m})}}$ (9)
@@
Given that there are $m^{2}$ nuggets, and, balance correlation considers 3 nuggets there are $m^{6}$ possible balance correlations. However these are not all unique, because transposition is a permutation and by definition A.2, nuggets include transposed matrices.
@@
Assume nuggets $\{U_{k},U_{l},U_{m}\}\in\mathfrak{U}$ and define $\{U_{k}=U_{q}^{{}^{\prime}},U_{l}=U_{r}^{{}^{\prime}},U_{m}=U_{s}^{{}^{\prime}}\}$ which implies $\{U_{q},U_{r},U_{s}\}\in\mathfrak{U}$. Since, transposing is a permutation, by equation 10 we have
@@
It follows that quantifying a sole statement requires a symmetric nugget $U_{k}$, which implies by lemma A.1 the matrices in $U_{k}$ are $M_{t}$ and $M_{t}^{{}^{\prime}}$, as well as that $U_{l}=U_{m}^{{}^{\prime}}$, which implies $U_{l}=M_{v}M_{w}^{{}^{\prime}}$ and $U_{m}=M_{w}M_{v}^{{}^{\prime}}$. In those cases, one and only one configuration of nuggets (and matrices) exists, since transposing equals the trivial identity permutation. If $U_{k}=U_{q}^{{}^{\prime}}$ for any $U_{k}=M_{t}\times M_{s}^{{}^{\prime}}$ and $U_{q}=M_{s}\times M_{t}^{{}^{\prime}}$ will contain the same set of values, but in a different order. Similarly, in any $U_{l}U_{m}$ that has $U_{s}U_{r}$ as a distinguishable transpose, all values $U_{l}U_{m}$ uniquely exist as transpose in $U_{s}U_{r}$, and vice verse. So for any $U_{l}=M_{v}M_{w}^{{}^{\prime}}$ and $U_{m}=M_{x}M_{y}^{{}^{\prime}}$ there exists $U_{r}=M_{w}M_{v}^{{}^{\prime}}$ and $U_{s}=M_{y}M_{x}^{{}^{\prime}}$, such that $\sum_{k}(M_{v_{ik}}M_{w_{ki}})(M_{x_{kj}}M_{y_{jk}})=\sum_{k}(M_{y_{jk}}M_{x_{kj}})(M_{v_{ik}}M_{w_{ki}})\,\forall\,\{i,j,k\}$. Hence, only when either or both variables in $\rho$ have a distinguishable transpose a distinguishable twin statement exists that produces the exact same value for $\rho$, i.e. showing duality for this subset of balance correlations.
@@
Definition A.3 determines any balance correlation $\rho(U_{k},U_{l}U_{m})$ is a bi-variate Pearson correlation. This leads us to consider two conditions, or distinguishing assumptions, as special cases. Although we recognise more complex situations could apply, these two limiting conditions will provide an upper and lower cardinality bound for $\mathfrak{B}$.
Assumption A.1.
@@
This implies that for all nuggets in =$\mathfrak{Z}=\{M_{r}\times M_{s}^{{}^{\prime}}\,|\,r\neq s,M_{r}\in\mathfrak{M},M_{s}^{{}^{\prime}}\in\mathfrak{M}^{{}^{\prime}}\}$ the variances $var(Z_{h})=0$. Since the Pearson correlation is defined if and only if $var(U_{k})>0$ and $var(U_{l}U_{m})>0$, the balance correlations that include $Z_{h}$ are not defined, and cannot be considered. Therefore, under assumption A.1 the nugget set cardinality will be $|\mathfrak{U}|=m$, and will only contain symmetric nuggets.
@@
Assumption A.2.
All matrices in $\mathfrak{M}$ are asymmetric, nuggets have $var(U_{k})>0\,\forall\,k\in\{1,...,g\}$, and the inner product of nuggets have $var(U_{l}U_{m})>0\,\forall\,l,m\in\{1,...,g\}$.
The proof of theorem 2.1 provides the conditions under which a specific configuration of nuggets doesn’t have a twin. It requires a symmetric nugget in the first variable in $\rho$, and a symmetric inner product of nuggets in the second variable. The latter condition can only hold when no more than two different matrices constitute the two nuggets. The order of the two matrices do define the specific nuggets, and hence is important. However, these two nuggets also need to be transposes of each other. Given $m$ asymmetric matrices there are exactly $m$ symmetric nuggets, according to the proof of lemma A.1. Furthermore, there are exactly $g=m^{2}$ nuggets, which includes their transposes. So, the total number of possible inner products equals $m^{4}$, of which $m^{2}$ will be inner products with the transpose. Hence, there are $\alpha=m\cdot m^{2}=m^{3}$ number of configurations of matrices in the variables of $\rho$ that quantify a sole statement. The other possible balance correlations all have exactly one other equal quantification of a twin statement. Therefore,
@@
where $m^{6}$ is the number of combinations of all possible nuggets (2 matrices in a nugget) in the three nuggets used in $\rho(U_{k},U_{l}U_{m})$. For any set of mixed symmetric and asymmetric matrices the cardinality of $|\mathfrak{B}|$ will be between those in equations 12 and 13.

### v2

@@ 4 Extensions
This paper deals with signed graphs, where edges between nation states are undirected symmetric ties. One of major advances in [27] was to recognize that networks are often, perhaps most often, composed of directed ties. The work presented here can be extended to directed ties, also, leading to new specific theoretical predictions and explorations of possible balance statements and commensurate balance correlations to assess them.
@@
However, it should be noted that this increases substantially the number of possible balance statements. Instead of one relation for each pair, there is both an in-tie and an out-tie that would have to be specified; thus, for each pair we must specify two relations, not just one. For example, we could describe a case where Ego sent a positive tie to Alter, but Alter does not respond with any sentiment (i.e., send a z-tie back). Perhaps Ego sends a negative tie to X, and X reciprocates with a non-tie. Further, suppose Alter sends a negative tie to X and X responds with a similar negative tie back to Alter. This configuration would be described as <pz . nz nn>. Since there are six directed ties in any labeled triple, then the number of possible directed triad signatures is $7293^{6}=$ 729, not simply $27$ as described in this paper. An interesting extension of our approach to digraphs is that of these $729$ signatures, $702$ have an indistinguishable twin and $27$ do not (see the proof of the theorem 2.1 in the appendix), resulting in $378$ unique balance correlations. We expound on these structures, their balance statements, and their empirical results in a subsequent paper on balance correlations in directed graphs.
An extension that would even further reintegrate a structural approach with Heider’s original ideas is to replace an individual node as context with sets of attitudes, values, norms, or issues [37]. This would allow to test different versions of balance theory, and further establish the link to cognitive dissonance theory [17, 23, 38].
The null model we have chosen controls for both positive and negative degree heterogeneity of the nodes. The result of [29] is very general, and the exponential family framework can be used to control for other social forces, such as geographic proximity and language, by augmenting or replacing the sufficient statistic in Eq 9 with operationalizations of these effects.
5 Discussion and conclusion
@@
Our intent here is to propose a new approach to the theorizing and study of balance processes in network contexts. By formally incorporating neutral or zero ties into the analysis, we extend balance theory in a significant way that better matches the real world of actors in a systemic relational structure. And, in an important way, we return to Heider’s original thinking by focusing on Ego’s choice of how to relate to a specific Alter when confronted with the concomitant relations with one or many third parties. We present a novel nomenclature, the triad signature, that assists in identifying and theorizing about (un)balance conditions.
@@
Finally, we propose a simple, intuitive measure, the balance correlation, that directly reflects the extent to which Heider’s predictions (and its extensions in this paper) are supported by the evidence in an observed network. This measure captures the difference in probabilities of Ego’s choice of a tie with Alter, conditioned on whether the specific balance-theoretic 2-path configuration exists or does not exist. We also describe a reasonable null test of this statistical measure and apply it to a well-known data set employed by political scientists in international affairs, the Correlates of War data.
@@
Second, we find that the predictions around unbalanced states are not confirmed for the most part and occasionally even counter to Heider’s original contention. For example, the theoretically unbalanced <n.pp> correlation hovers around zero for most of the two centuries of data. And a clear example of a result that dis-confirms an unbalanced prediction is given in Fig 7, where the theoretically unbalanced <p.pn> signature is consistently positive in the post WWII era. This is consistent with an alternative to balance seeking behavior, namely a tertius gaudens strategy as described in [5]. However, the results also show that the balanced <n.pn> signature is stronger. This emphasizes that both strategies may be at play at the same time in the same network, revealing the complexity of studying international relations.
Thirdly, in a related general finding, it appears that the z-ties (neutral ties) play a larger role in establishing balance, or at least avoiding unbalanced states, than negative hostile ties over substantial time periods. For example, in the case of where the Ego  Alter dyad is embedded in many <pp> 2-path relations (see Fig 6), there is no discernible evidence that Ego avoids negative ties (<n.pp>), but there is substantial and significant suppression or avoidance of neutral ties (<z.pp>).
This study of embedded triadic balance behaviour over time suggests that although the strive for balance may be a component of individual countries’ motivation, it does not necessarily suggests social equilibrium or social resilience of the system. In fact, we find over time that balance can become more and less prevalent in whole networks and that these levels of prevalence can be stable over longer time spans. This is consistent with Heider’s [26] observation that the tension produced by unbalance may be attractive. In the setting of international relations such imbalance may hold promise for opportunities to achieve nodal benefits. The presented method indicates the prevalence of different types of balance, imbalance and undetermined behaviors.
In sum, it is our contention that balance correlations provide a simple, intuitive way to assess the extent to which specific assumption on balance behaviors might be true on the average across an entire network. Furthermore, by incorporating agentic zeroes into the exploration of balance conditions, we may uncover new properties of social order that have been prevalent, even dominant, but hidden from the researcher’s purvue or agenda. We hope this new approach opens up avenues of exploration that have been heretofore ignored.
As a final remark we want to emphasize that an updated Correlates of War database (real-time) may demonstrate the practical use of balance correlation as an indicator of the level of global stability. Although we leave it to historians to corroborate that the LOA-era was characterized by dismal turbulence more than in other periods, the consistent drop in balanced relations might well serve as warning signal. If so, balance correlations may inform politicians and military leaders on global temperature in international relations.
Supporting information
S1 Appendix.
Proof of theorem 2.1 The claim in Theorem 2.1 is based on the fact that if the indices of the two variables in the balance correlation are permuted in the same way, the correlation does not change due to associative and commutative properties of the addition of the cross-products in the correlation function. Transposing matrices is one unique permutation of the elements in a matrix.
Definition 5.1.
@@
Proof of lemma 5.1.
Note that $M_{q_{ji}}=M_{q_{ij}}^{{}^{\prime}}$. By contradiction, assume for any $ij$-th element in a nugget, the element-wise product of matrices, $M_{r_{ij}}\times M_{s_{ji}}=1$ and $M_{r_{ji}}\times M_{s_{ij}}=1$. Since, $M_{q_{ij}}\in\{0,1\}$, this implies $M_{r_{ij}}=M_{s_{ji}}=1$ and $M_{r_{ji}}=M_{s_{ij}}=1$, so $M_{r_{ij}}+M_{s_{ij}}=2$. However, if $r\neq s$ this contradicts that the sum over all matrices results in a matrix $\mathfrak{L}$ (see definition 5.1). If and only if $r=s$ the assumption holds, as the summation visits each matrix $q$ in $\mathfrak{M}$ only once, which implies there are exactly $m$ symmetric nuggets; those of a matrix and it’s transpose. ∎
Definition 5.3.
@@
$\rho(U_{k},U_{l}U_{m})=\frac{\operatorname{cov}(U_{k},U_{l}U_{m})}{\sqrt{\operatorname{var}(U_{k})\operatorname{var}(U_{l}U_{m})}}$ (10)
@@
Given that there are $m^{2}$ nuggets, and that balance correlation considers 3 nuggets, there are $m^{6}$ possible balance correlations. However these are not all unique, because transposition is a permutation and by definition 5.2, nuggets include transposed matrices.
@@
Assume nuggets $\{U_{k},U_{l},U_{m}\}\in\mathfrak{U}$ and define $\{U_{k}=U_{q}^{{}^{\prime}},U_{l}=U_{r}^{{}^{\prime}},U_{m}=U_{s}^{{}^{\prime}}\}$ which implies $\{U_{q},U_{r},U_{s}\}\in\mathfrak{U}$. Since, transposing is a permutation, by Eq 11 we have
@@
It follows that quantifying a sole statement requires a symmetric nugget $U_{k}$, which implies by lemma 5.1 the matrices in $U_{k}$ are $M_{t}$ and $M_{t}^{{}^{\prime}}$, as well as that $U_{l}=U_{m}^{{}^{\prime}}$, which implies $U_{l}=M_{v}M_{w}^{{}^{\prime}}$ and $U_{m}=M_{w}M_{v}^{{}^{\prime}}$. In those cases, one and only one configuration of nuggets (and matrices) exists, since transposing equals the trivial identity permutation. If $U_{k}=U_{q}^{{}^{\prime}}$ for any $U_{k}=M_{t}\times M_{s}^{{}^{\prime}}$ and $U_{q}=M_{s}\times M_{t}^{{}^{\prime}}$ will contain the same set of values, but in a different order. Similarly, in any $U_{l}U_{m}$ that has $U_{s}U_{r}$ as a distinguishable transpose, all values $U_{l}U_{m}$ uniquely exist as transpose in $U_{s}U_{r}$, and vice verse. So for any $U_{l}=M_{v}M_{w}^{{}^{\prime}}$ and $U_{m}=M_{x}M_{y}^{{}^{\prime}}$ there exists $U_{r}=M_{w}M_{v}^{{}^{\prime}}$ and $U_{s}=M_{y}M_{x}^{{}^{\prime}}$, such that $\sum_{k}(M_{v_{ik}}M_{w_{ki}})(M_{x_{kj}}M_{y_{jk}})=\sum_{k}(M_{y_{jk}}M_{x_{kj}})(M_{v_{ik}}M_{w_{ki}})\,\forall\,\{i,j,k\}$. Hence, only when either or both variables in $\rho$ have a distinguishable transpose a distinguishable twin statement exists that produces the exact same value for $\rho$, i.e. showing duality for this subset of balance correlations.
@@
Definition 5.3 determines any balance correlation $\rho(U_{k},U_{l}U_{m})$ is a bi-variate Pearson correlation. This leads us to consider two conditions, or distinguishing assumptions, as special cases. Although we recognise more complex situations could apply, these two limiting conditions will provide an upper and lower cardinality bound for $\mathfrak{B}$.
Assumption 5.1.
@@
This implies that for all nuggets in =$\mathfrak{Z}=\{M_{r}\times M_{s}^{{}^{\prime}}\,|\,r\neq s,M_{r}\in\mathfrak{M},M_{s}^{{}^{\prime}}\in\mathfrak{M}^{{}^{\prime}}\}$ the variances $\operatorname{var}(Z_{h})=0$. Since the Pearson correlation is defined if and only if $\operatorname{var}(U_{k})>0$ and $\operatorname{var}(U_{l}U_{m})>0$, the balance correlations that include $Z_{h}$ are not defined, and cannot be considered. Therefore, under assumption 5.1 the nugget set cardinality will be $|\mathfrak{U}|=m$, and will only contain symmetric nuggets.
@@
Assumption 5.2.
All matrices in $\mathfrak{M}$ are asymmetric, nuggets have $\operatorname{var}(U_{k})>0\,\forall\,k\in\{1,...,g\}$, and the inner product of nuggets have $\operatorname{var}(U_{l}U_{m})>0\,\forall\,l,m\in\{1,...,g\}$.
The proof of theorem 2.1 provides the conditions under which a specific configuration of nuggets doesn’t have a twin. It requires a symmetric nugget in the first variable in $\rho$, and a symmetric inner product of nuggets in the second variable. The latter condition can only hold when no more than two different matrices constitute the two nuggets. The order of the two matrices do define the specific nuggets, and hence is important. However, these two nuggets also need to be transposes of each other. Given $m$ asymmetric matrices there are exactly $m$ symmetric nuggets, according to the proof of lemma 5.1. Furthermore, there are exactly $g=m^{2}$ nuggets, which includes their transposes. So, the total number of possible inner products equals $m^{4}$, of which $m^{2}$ will be inner products with the transpose. Hence, there are $\alpha=m\cdot m^{2}=m^{3}$ number of configurations of matrices in the variables of $\rho$ that quantify a sole statement. The other possible balance correlations all have exactly one other equal quantification of a twin statement. Therefore,
@@
where $m^{6}$ is the number of combinations of all possible nuggets (2 matrices in a nugget) in the three nuggets used in $\rho(U_{k},U_{l}U_{m})$. For any set of mixed symmetric and asymmetric matrices the cardinality of $|\mathfrak{B}|$ will be between those in Eq 13 and Eq 14.
@@
Acknowledgments
We like to thank participants at Networks 2021 (Bloomington, USA), Sunbelt 2022, (Cairns, Australia), Sunbelt 2023 (Portland, USA), ION IX (Lexington, USA), and EUSN 2023 (Ljubljana, Slovenia) for valuable remarks.